The volume of water entering a tank from a pump is related by the formula v=ah^3+bh^2+ch+d where is the height of the tank, a, b, c and d are constants.
Fine the condition for minimum and maximum water storage.
1
Expert's answer
2020-11-25T16:22:31-0500

"V=ah^3+bh+ch+d"

at Maximum and Minimum Volume, "\frac{dV}{dh}=0"

getting the derivative of V w.r.t to h we obtain

"\frac{dV}{dh}=3ah^2+2bh+c"

this implies that, "3ah^2+2bh+c=0"

solving for h using that maximizes/minimizes Volume using completing square we get:

"3ah^2+2bh=-c"

dividing by 3a, "h^2+\frac{2b}{3a}h=\frac{-c}{3a}"

completing the square, "h^2+\frac{2b}{3a}h+(\frac{b}{3a})^2=\frac{-c}{3a}+\frac{b^2}{9a^2}"

"(h+\frac{b}{3a})^2= \frac{b^2-3ac}{9a^2}"

"(h+\frac{b}{3a})=\sqrt\frac{b^2-3ac}{9a^2}"

"h1= \frac{b}{3a}+\frac{\sqrt {(b^2-3ac)}}{3a}"

"h1= \frac{-b+\sqrt{(b^2-3ac)}}{3a}"

"h2= \frac{b}{3a}-\frac{\sqrt{(b^2-3ac)}}{3a}"

"h2= \frac{-b-\sqrt{(b^2-3ac)}}{3a}"


therefore, the volume will be maximum/minimum when


"h1= \frac{-b+\sqrt{(b^2-3ac)}}{3a}" and


"h2= \frac{-b-\sqrt{(b^2-3ac)}}{3a}" (condition)


At Maximum

"\frac{dV}{dh}<0" at H1 and H2


At minimum

"\frac{dV}{dh}>0" at H1 and H2